Could you please help me with this question. It came from a study guide given in class. Thank you in advance!
Chapter Lesson: Confidence Intervals for the Mean, Small Samples (n<30)
6. Find the maximum error of estimate (E), and the small sample confidence interval for the population mean for the given values of c, x, σ, and n.
c = 95%
x = 162
σ = 10
n = 16

Answer by stanbon(75887)

Find the maximum error of estimate (E), and the small sample confidence interval for the population mean for the given values of c, x, σ, and n.
c = 95%
x = 162
σ = 10
n = 16
----
E = t
------
t = inv(0.975,15) = 2.13
----
E = 2.13*10/sqrt(16) = 5.325
================
Cheers,
Stan H.
==============